OpenStudy (anonymous):

Complete the table for the radioactive isotope. (Round to 2 decimal places.) Isotope - ^226 Ra Half-life (years)- 159 Initial quanity ??? g (unknown) Amount after 1000 years - 1.95g

OpenStudy (anonymous):

\[Q_0\times \left(\frac{1}{2}\right)^{\frac{1000}{159}}=1.95\]solve for \(Q_0\)

OpenStudy (anonymous):

by division and a calculator

OpenStudy (anonymous):

\[Q_0=1.95\div\left(\frac{1}{2}\right)^{\frac{1000}{159}}\]